We are given the following information about the assets and liabilities of a bank:Liabilities $65 million Deposits Assets Loans Required Reserves Excess Reserves $2 million Treasury Securities $5 million

a. The Fed sets a reserve requirement of 3% on deposits between $16 million and $122 million. If the bank holds $5 million dollars in US Treasury Securities and $2 million in excess reserves, compute the bank’s required reserve level and the quantity of loans this bank is able to make to the public. b. What is the value of the money multiplier? [Money Multiplier = 1/Effective Reserve Ratio; Effective Reserve Ratio = (Required Reserves + Excess Reserves)/Deposits] c. Suppose that the Fed increases the required reserve ratio to 5% on deposits between $16 million and $122 million. Further, the bank expands its holdings of US Treasury Securities to $6.5 million and holds $3 million in excess reserves. Calculate the bank’s new required reserve level and the quantity of new loans which this bank is able to make to the public, assuming no change in the quantity of deposits. d. What is the new value of the money multiplier? Is the change in the reserve requirement an example of expansionary or contractionary monetary policy?
Answer #1

Hi,

Hope you are doing well!

Question:

Answer:

a). Answer:

Deposits = $65 million

Reserve requirement is = 3%

[Fed sets a reserve requirement of 3% on deposits between $16 million and $122 million]

Bank’s required reserve level = $65 million * 3% = $1.95 million

Loanable Fund of the bank = $65 million - $1.95 million = $63.05 million

b). Answer:

Money Multiplier = 1/Effective Reserve Ratio

1/3% = 1/0.03 = 33.33

Money Multiplier = 33.33

Effective Reserve Ratio = (Required Reserves + Excess Reserves)/Deposits]

= ($1.95 million + $2 million)/ $65 million

= 3.95/65 = 0.06076 = 6.07

Effective Reserve Ratio = 6.07

c). Answer:

Fed increases the required reserve ratio to 5% on deposits between $16 million and $122 million.

The bank expands its holdings of US Treasury Securities to $6.5 million and holds $3 million in excess reserves.

Assuming no change in the quantity of deposits.

Required reserve level = $65 million * 5% = $3.25 million

The quantity of new loans which this bank is able to make to the public = $65 million - $3.25 million = $61.75 million

The quantity of new loans which this bank is able to make to the public = $61.75 million

d). Answer:

The change in the reserve requirement from 3% to 5% is an example of contractionary monetary policy.

Contractionary monetary policy: It is type of monetary policy in which the central bank reduce the money supply through increasing discount rate or reserve requirement or selling government securities.

Thank You